I AM HAVING PROBLEM IN SOLVING THIS QUESTION KINDLY HELP ME............
EACH INCH ON RULER A IS MARKED IN EQUAL 1/8 INCH UNITS,AND EACH INCH ON RULER B IS MARKED IN EQUAL 1/12 INCH UNITS.WHEN RULER A IS USED,A SIDE OF A TRIANGLE MEASURES 12 OF THE 1/8 INCH UNITS.WHEN RULER B IS USED,HOW MANY 1/12 INCH UNITS WILL THE SAME SIDE MEASURE?
(PLEASE EXPLAIN ME TOO WHAT ARE THEY ASKING ABOUT AND HOW YOU SOLVED IT)

Answer by Theo(13342)

ruler A is marked in 1/8 of an inch units.
ruler B is marked in 1/12 of an inch units.
with ruler A, the side of the triangle measures 12 of the 1/8 inch units.
this means that the side of the triangle is equal to 12/8 inches in length.
you essentially want to convert 12/8 into something over 12.
let that something be represented by x.
your equation becomes:
12/8 = x/12
cross multiply to get 8x = 12*12 which becomes:
8x = 144
divide both sides of this equation by 8 to get:
x = 144/8 which becomes:
x = 18
your solution is that 12/8 is equivalent to 18/12
if you simplify 12/8, it becomes 3/2.
if you simplify 18/12, it becomes 3/2.
the two measures are equivalent.